OpenStudy (anonymous):

How do you integrate (1+x)^-2

myininaya (myininaya):

let u=1+x => du=dx \[\int\limits_{}^{}u^{-2} du =\frac{u^{-2+1}}{-2+1}+C\] don't forget to replace u with 1+x

OpenStudy (anonymous):

thanks